OpenStudy (anonymous):

For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis? negative 3 over 2 negative 2 over 3 2 3

OpenStudy (dumbcow):

ok you need to factor this or use quasratic formula

OpenStudy (dumbcow):

*quadratic

OpenStudy (anonymous):

2

OpenStudy (anonymous):

It will cross x when y = 0. Assuming you mean y = 2x^2 - 7x +6 x crosses when 2x^2 - 7x +6 = 0 2x^2 - 7x +6 = (2x - 3)(x - 2) Therefore, y = 0 when x = 2 and x = 3/2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

oh- my calc missed the 3/2